A linear regression equation has b = 3 and a = – 6. What is the predicted value of Y for X = 4?

a. ​Ŷ = 6

b. ​Ŷ = –21

c. ​Cannot be determined without additional information

d. ​Ŷ = –2
Answer #1

soli 6:3 Q:-6 X=4 We know that ŷ = a +bx - -6+3(+) y=-6+12 Û = 6 A. . optional is correct